A boat travels upstream for 3 hours. The return trip requires 2 hours. If the speed of the current is 2 km per hour, find the speed of the boat in still water.
Answer by mananth(16946)

Boat speed =x kph
current speed 2 kph

time with current 2 hours
time against current 3 hours
Distance = same
8(x+2 ) = 3 ( x -2 )
8 x 16 = 3 x -6
3 x -8 x = 16 -6
x = 22 kph Boat speed